OpenStudy (anonymous):

three pipes can fill a tank.the diameters of the pipes are 1cm,1 1/3 cm and 2cm respectively.at a certain time the amount of water flowing from each pipe is proportionate to the square of their diameters.the pipe with the largest diameter can fill the tank in 61 minutes.then in how much time will all the three pipes fill the tank if they are opened together?

OpenStudy (whpalmer4):

Ah, a fun variation on a common problem :-) First step is to figure out the rates of flow for each of the 3 pipes. Can you do that?

OpenStudy (anonymous):

actually i got the solution see we know first and the largest pipe can fill alone in 61 mins taking t a variable t/61∝2*2 t/61=4k k=t/61*4 similarly putting value of k for other two pipes k=16t/61*4*9 and k=4t/61*4 adding all values of k we get t=36

OpenStudy (whpalmer4):

Yes, that's the correct answer, good job!